### v1

# Geometric Height on flag varieties in positive characteristic

Abstract.
Let $k$ be an algebraically closed field of characteristic $p\neq 0$. Let $G$ be a connected reductive group over $k$, $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow G$ be a strictly anti-dominant character. Let $C$ be a projective smooth curve over $k$ with function field $K=k(C)$ and $F$ be a principal $G$-bundle on $C$. Then $F/P\longrightarrow C$ is a flag bundle and $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ on $F/P$ is a relatively ample line bundle. We compute the height filtration and successive minima of the height function $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ over the flag variety $X=(F/P)_{K}$.
1. Introduction
1.1. Height filtration and successive minima
Let $K$ be either a number field or $K=k(C)$ where $C$ is a projective smooth curve over a field $k$. Let $X$ be a projective variety of dimension $d$ over $K$ and $\overline{L}$ be an adelic line bundle on $X$. These data induce an Arakelov height function $h_{\overline{L}}$ on $X$ (see [9, §9] for a survey). A typical case is the geometric height, which is the one we concern in this article. Here we give the definition.
If $K=k(C)$, consider a projective flat morphism $\mathcal{X}\longrightarrow C$ with the generic fiber $X\longrightarrow\operatorname{Spec}(K)$ and a line bundle $\mathcal{L}$ on $\mathcal{X}$ with $\mathcal{L}_{K}\simeq L$. The data $(\mathcal{X},\mathcal{L})$ define an adelic line bundle $\overline{L}$ and the height function $h_{\overline{L}}$ is given by
$h_{\overline{L}}:X(\overline{K})\longrightarrow\mathbb{Q},\;x\longmapsto\frac{\mathcal{L}\cdot\overline{\{x\}}}{{\mathrm{deg}}(x)}\;\;\text{where $\overline{\{x\}}$ is the closure of $x$ in $\mathcal{X}$}.$
We also denote this by $h_{\mathcal{L}}$ if there is no ambiguity. If $K$ is a number field, the height function can be defined similarly via arithmetic intersection theory.
For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\overline{L}}(x)<t\big{\}}$. Let’s call $\big{\{}Z_{t}:t\in\mathbb{R}\big{\}}$ the height filtration, and call its jumping points the successive minima. Note that our definition of successive minima are slightly different with Zhang [10]. Zhang considers only dimension jumps $e_{i}=\inf\big{\{}t:\dim Z_{t}\geq d-i+1\big{\}}$, $i=1,\dots,d+1$. In this article, $e_{i}$ will be called the successive minima of Zhang to avoid ambiguity.
In [3], Fan-Luo-Qu provides a new case where height filtration can be explicitly computed, which is the geometric height on flag varieties under the assumption ${\rm char}(k)=0$. Following their essence, we examine the case of positive characteristic.
Let $k$ be an algebraically closed field and $C$ be a projective smooth curve over $k$ with function field $K=k(C)$. Let $G$ be a connected reductive group over $k$, $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly antidominant character. Let $F$ be a principal $G$-bundle over $C$ and $\mathcal{X}=F/P$ with generic fiber $X=(F/P)_{K}$. Let $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ and $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ be the induced height function by the following diagram.
${X=(F/P)_{K}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{X}=F/P\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Spec}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C}$
Let $F_{Q}$ be the canonical reduction of $F$ to $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the induced cocharacter. Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$. For $w\in W_{Q}\backslash W/W_{P}$, let $C_{w}=(F_{Q}\times_{Q}QwP/P)_{K}\subseteq X$ be the corresponding Schubert cell. The numbers $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ are well-defined. One main result in [3] is the following.
Theorem 1.1 (Theorem 2.1, [3]).
Assume $k$ has characteristic zero. For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\mathcal{L}_{\lambda}}(x)<t\big{\}}$. Then
$Z_{t}=X\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}C_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t}C_{w}.$
From now on assume $k$ has positive characteristic unless other description is given. We show that this result does not hold in characteristic $p$ unless one puts the following assumption on $F$, namely $F$ admits a strongly canonical reduction.
Definition 1.2 (Definition 2.2).
We say a reduction $F_{Q}$ to a parabolic subgroup $Q$ is strongly canonical if for any non-constant finite morphism $f:C^{\prime}\to C$, the pullback $f^{*}(F_{Q})$ is the canonical reduction of $f^{*}F$.
Theorem 1.3 (Theorem 3.1).
The result in Theorem 1.1 holds under the assumption that $F$ admit a strongly canonical reduction.
Canonical reduction is automatically strongly canonical in characteristic 00, whereas strongly canonical reduction may not exist in characteristic $p>0$. For general $G$-bundles that may not admit a strongly canonical reduction, we have the following rough form of our another main result.
Theorem 1.4.
The height filtration of $X$ is given by successively deleting some Frobenius twist of Schubert cells.
More explicitly, a theorem by Langer [5] shows for a principal $G$-bundle $F$ the strongly canonical reduction exists for $({\rm Fr}^{n})^{*}F$, when $n$ is sufficiently large. Therefore, consider the following Cartesian diagram for $n$ sufficiently large.
${(\operatorname{Fr_{C}}^{n})^{*}F/P}$${F/P}$${C}$${C}$$\scriptstyle{\phi}$$\scriptstyle{\operatorname{Fr}_{C}^{n}}$
with its generic fiber
[…]
Here $\operatorname{Fr}_{C}$ is the absolute Frobenius on $C$. Suppose $n$ is large enough such that $(\operatorname{Fr}_{C}^{n})^{*}F$ has strongly canonical reduction, then we have:
Theorem 1.5 (Theorem 3.2).
The height filtration of $h_{\mathcal{L}_{\lambda}}$ on $X$ is given by the image of the height filtration of the height filtration of $\tilde{X}$ along the homeomorphism $\phi_{K}$, the successive minima of $X$ are $\frac{1}{p^{n}}$ of the succesive minima of $\tilde{X}$.
Remark 1.6.
Here the height filtration of $\tilde{X}$ is computed by Theorem 3.1. Thus, we say the height filtration of $X$ is given by successively deleting some Frobenius twist of Schubert cells as stated in Theorem 1.4.
1.2. Toy example: projective spaces
Let $k$ be an algebraically closed field of positive characteristic and $C$ be a curve over $k$ with function field $K=k(C)$. Let $E$ be a vector bundle of rank $n$ on $C$. We take $\mathcal{X}=\mathbb{P}(E)$, $\mathcal{L}$ to be the relative ample line bundle $\mathcal{O}_{\mathbb{P}(E)}(1)$ and $X=\mathbb{P}(V)$ where $V$ is the generic fiber of $E$. A special case of a theorem by Ballaÿ [1] show that: Let $\zeta_{ess}:=\inf_{t\in\mathbb{R}}\{Z_{t}=X\}$. Then
$\zeta_{ess}=\lim_{n}\frac{\mu_{max}(\pi_{*}\mathcal{L}^{\otimes n})}{n}(=\lim_{n}\frac{\mu_{max}({\rm Sym}^{n}E)}{n}).$
where $\mu_{\max}(-)$ is the maximal slope appears in the Harder-Narasimhan filtration.
Warning 1.7.
In positive characteristic, the symmetric power of a semistable vector bundle may not remain semistable, making the left-hand side quite difficult to compute. It’s important to note that this issue arises only in positive characteristic. However, this problem can be resolved by sufficiently twisting the Frobenius morphism multiple times.
Proposition 1.8 (Ramanan, Ramanathan, [7]).
For above $E$ semistable, there exists $N$ large enough such that for $n\geq N$, any symmetric power of ${\rm Fr}^{n,*}_{C}E$ is semistable, where ${\rm Fr}_{C}$ is the absolute Frobenius of $C$.
Theorem 1.9 (Baby-version).
Fix a sufficiently large $n$ in the sense of the above proposition. Denote $V^{\prime}={\rm Fr}_{K}^{n,*}V$ and $\{V^{\prime}_{i}\}$ the generic fiber of Harder-Narasimhan filtration $\{E^{\prime}_{i}\}$ of ${\rm Fr}_{C}^{n,*}E$.
  1. (1)
    The height filtration of $\mathbb{P}(V^{\prime})$ is given by $\mathbb{P}(V^{\prime})\supsetneq\mathbb{P}(V^{\prime}/V^{\prime}_{1})\supsetneq\mathbb{P}(V^{\prime}/V^{\prime}_{2})\cdots\supsetneq\mathbb{P}(V^{\prime}/V^{\prime}_{r})=\emptyset$ with successive minima $\mu_{\rm max}(E^{\prime}_{i}/E^{\prime}_{i-1})$.
  2. (2)
    The height filtration of $h_{\mathcal{O}(1)}$ on $\mathbb{P}(V)$ is the image of the height filtration of $h_{\mathcal{O}(1)}$ on $\mathbb{P}(V^{\prime})$ along the natural map. The successive minima of $\mathbb{P}(V)$ are $\frac{1}{p^{n}}\mu_{\rm max}(E^{\prime}_{i}/E^{\prime}_{i-1})$.
A byproduct would be the following equality:
$L_{max}=\lim_{n}\frac{\mu_{max}({\rm Sym}^{n}E)}{n},$
where $L_{max}:=\max_{f:Y\to C}\{\mu_{max}(f^{*}E)\}$, where $f$ runs through all finite non-constant morphisms. It would be in particular interesting to have a purely vector bundle theoretical proof of this equality. Note that $L_{max}=\mu_{max}$ in characteristic zero, and $L_{max}=\max_{n}\{\mu_{max}(Fr_{C}^{n,*}E)\}$ in positive characteristic.
1.3. Acknowledgments
We sincerely thank Dr. Binggang Qu for his invaluable advice and support on this work. This research also stems from the 2024 Algebra and Number Theory Summer School held at Peking University. We are deeply grateful to the organizers for their efforts and hospitality.
2. Strongly canonical reduction
2.1. Vector bundles
Let $E$ be a vector bundle on $C$. The degree of $E$ is ${\mathrm{deg}}(E):={\mathrm{deg}}(\det(E))$ and the slope of $E$ is $\mu(E):={\mathrm{deg}}(E)/\operatorname{rk}(E)$. It is called slope semistable if for every subbundle $F$ of $E$, $\mu(F)\leq\mu(E)$. This is equivalent to $\mu(Q)\geq\mu(E)$ for every quotient bundle $Q$ of $E$.
There exists uniquely a filtration $0=E_{0}\subseteq E_{1}\subseteq\cdots\subseteq E_{r}=E$ such that
  1. (1)
    $E_{i}/E_{i-1}$ is semistable;
  2. (2)
    $\mu(E_{i}/E_{i-1})>\mu(E_{i+1}/E_{i})$.
This filtration is called the Harder-Narasimhan filtration of $E$. $\mu(E_{1})$ is called maximal slope of $E$ and is also denoted by $\mu_{\max}(E)$.
2.2. Principal bundles
For any linear algebraic group $\Gamma/k$, let $X(\Gamma)=\mathrm{Hom}(\Gamma,\mathbb{G}_{m})$ denote the character group of $\Gamma$. For a cocharacter $f\in X(\Gamma)^{\vee}$ and a character $\lambda\in X(\Gamma)$, we shall denote the pairing by $\langle f,\lambda\rangle$. For any $\lambda\in X(\Gamma)$, denote by $k_{\lambda}$ the one-dimensional representation on the vector space $k$ with $\Gamma$ acting by $\lambda$.
A principal $\Gamma$-bundle on $C$ is a variety $F$ equipped with a right action of $\Gamma$ and a $\Gamma$-equivariant smooth morphism $F\longrightarrow C$ such that the map
$F\times_{C}(C\times\Gamma)\longrightarrow F\times_{C}F,\quad(f,(x,g))\longmapsto(f,fg)$
is an isomorphism.
Attached to a principal $\Gamma$-bundle $F$, one has an associated cocharacter
${\mathrm{deg}}(F):X(\Gamma)\longrightarrow\mathbb{Z},\quad\lambda\longmapsto\langle{\mathrm{deg}}(F),\lambda\rangle={\mathrm{deg}}(F\times_{\Gamma}k_{\lambda}).$
Here ${\mathrm{deg}}(F\times_{\Gamma}k_{\lambda})$ is the degree of the line bundle $F\times_{\Gamma}k_{\lambda}$ on the curve $C$.
Let $H$ be a closed subgroup of a linear algebraic group $\Gamma$ over $k$. A reduction of structure group of $F$ to $H$ is a pair $(F_{H},\phi)$ where $F_{H}$ is a principal $H$-bundle and $\phi:F_{H}\times_{H}\Gamma\simeq F$ is an isomorphism.
By the universal property of the quotient $F/H$, the assignment to a section $\sigma:C\longrightarrow F/H$ the reduction $\sigma^{*}F$ of $F$ to $H$ is a one-one correspondence between reductions of structure group of $F$ to $H$ and sections of $F/H\longrightarrow C$.
2.3. Reductive groups, characters and cocharacters
Let $G$ be a connected reductive group over $k$. Fix a Borel subgroup $B\subseteq G$ and a maximal torus $T\subseteq B$. Let $W$ be the Weyl group and $\Delta$ be the set of simple roots with respect to $(G,B,T)$. For any $\alpha\in\Delta$, we denote by $\alpha^{\vee}$ the corresponding simple coroot.
We shall consider only parabolic subgroups containing $B$. For such a parabolic subgroup $P$, let $W_{P}\subseteq W$ be the Weyl group $W(L_{P})$ of the Levi factor $L_{P}\subseteq P$ and $\Delta_{P}\subseteq\Delta$ be the simple roots of $L_{P}$. Note that the natural inclusion
$X(P)\longrightarrow X(L_{P})\longrightarrow X(Z(L_{P}))$
becomes an isomorphism after tensoring with $\mathbb{Q}$. Thus we have
$X(T)_{\mathbb{Q}}\longrightarrow X(Z(L_{P}))_{\mathbb{Q}}=X(P)_{\mathbb{Q}}$
and by taking duals, we get the so-called slope map $X(P)^{\vee}_{\mathbb{Q}}\longrightarrow X(T)^{\vee}_{\mathbb{Q}}$ introduced in [8, §2.1.3]. In other words, a cocharacter on $X(P)_{\mathbb{Q}}$ can be extended canonically to $X(T)_{\mathbb{Q}}$.
2.4. The strongly canonical reduction to a parabolic subgroup
Let $G$ be a connected reductive group over $k$ and $F$ be a principal $G$-bundle over $C$. Let $F_{P}$ be a reduction of $F$ to a parabolic subgoup $P$. Let ${\mathrm{deg}}(F_{P})\in X(T)_{\mathbb{Q}}^{\vee}$ be the induced (rational) cocharacter.
The $G$-bundle $F$ is called semistable if for any parabolic subgroup $P$, any reduction $F_{P}$ of $F$ to $P$ and any dominant character $\lambda$ of $P$ which is trivial on $Z(G)$, we have $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle\leq 0$.
Definition 2.1.
We say $F$ is strongly semistable if for any non-constant finite morphism $f:C^{\prime}\to C$, the pullback $f^{*}F$ is semistable.
Among all filtrations of a vector bundle, there is a canonical one (the Harder-Narasimhan filtration). Similarly, among all reduction to parabolic subgroups, there is a canonical one. A reduction $F_{P}$ of $F$ to a parabolic subgroup $P$ is called canonical if the following two conditions hold:
  1. (1)
    The the principal $L_{P}$ bundle $F_{P}\times_{P}L_{P}$ is semistable.
  2. (2)
    For any non-trivial character $\lambda$ of $P$ which is non-negative linear combination of simple roots, $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle>0$.
Definition 2.2.
We say $F_{Q}$ is strongly canonical if for any non-constant finite morphism $f:C^{\prime}\to C$, the pullback $f^{*}(F_{Q})$ is canonical.
Remark 2.3.
If the base field has characteristic 00, then semistability (resp. canonical reduction) and strongly semistability (resp. strongly canonical reduction) are equivalent and the (strongly) canonical reduction exists uniquely [2]. However, in positive characteristic, the notion of strongly canonical reduction becomes essential for our purposes, despite the fact that it does not generally exist.
3. Height filtrations and successive minima
Let $C$ be a curve over a field $k$ of positive characteristic and $K$ be its function field. Let $G$ be a connected reductive group over $k$. Let $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly antidominant character. Let $F$ be a principal $G$-bundle over $C$. Let $\mathcal{X}=F/P$ and $X=(F/P)_{K}$. Let $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ and $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ be the induced height function.
${X=(F/P)_{K}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{X}=F/P\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Spek}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C}$
Let $F_{Q}$ be the canonical reduction of $F$ to $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the degree cocharacter. Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$. For $w\in W_{Q}\backslash W/W_{P}$, let $C_{w}=(F_{Q}\times_{Q}QwP/P)_{K}\subseteq X$ be the corresponding Schubert cell. We have the following theorem.
Theorem 3.1.
Assume $F$ admits strongly canonical reduction. For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\mathcal{L}_{\lambda}}(x)<t\big{\}}$. Then
$Z_{t}=X\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}C_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t}C_{w}.$
In particular, successive minima are $\zeta_{w}=\left\langle\operatorname{deg}\left(F_{Q}\right),w\lambda\right\rangle$ and Zhang’s successive minima are $e_{i}=\min\left\{\zeta_{w}:\ell(w)=\operatorname{dim}G/P-i+1\right\}$ where
$\ell(w)=\max_{\sigma\in W_{Q}wW_{P}}\min_{\tau\in\sigma W_{P}}\ell(\tau).$
For general $G$-bundle $F$ which does not necessarily admits a strongly canonical reduction, consider the following Cartesian diagram for $n$ sufficiently large.
${(\operatorname{Fr_{C}}^{n})^{*}F/P}$${F/P}$${C}$${C}$$\scriptstyle{\phi}$$\scriptstyle{\operatorname{Fr}_{C}^{n}}$
with its generic fiber
[…]
Theorem 3.2.
The height filtration of $h_{\mathcal{L}_{\lambda}}$ on $X$ is given by the image of the height filtration of the height filtration of $\tilde{X}$, the successive minima are $\frac{1}{p^{n}}$ of the successive minima of $\tilde{X}$.
3.1. A height lower bound in Schubert cells
In subsection 3.1 we follow [3] since there is no difference between characteristic zero and characteristic $p$. For $w\in W_{Q}\backslash W/W_{P}$, write $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P,\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P,C_{w}=\mathcal{C}_{w,K}$ and $X_{w}=\mathcal{X}_{w,K}$.
Let $F_{Q}$ be the canonical reduction of $F$ to $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the degree cocharacter. Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$. For $w\in W_{Q}\backslash W/W_{P}$, write $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P$, $\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P$, $C_{w}=\mathcal{C}_{w,K}$ and $X_{w}=\mathcal{X}_{w,K}$.
Note that for any $w^{\prime}\in W_{Q}$ and $\lambda\in X(T)$, we have $w^{\prime}\lambda-\lambda\in\mathbb{Z}\left[\Delta_{Q}\right]$ and consequently $\left\langle\operatorname{deg}\left(F_{Q}\right),w^{\prime}\lambda\right\rangle=\left\langle\operatorname{deg}\left(F_{Q}\right),\lambda\right\rangle$. Note also that for any $\lambda\in X(P)$ and $w\in W_{P},w\lambda=\lambda$. So the number $\left\langle\operatorname{deg}\left(F_{Q}\right),w\lambda\right\rangle$ is well-defined for any $\lambda\in X(P)$ and $w\in W_{Q}\backslash W/W_{P}$.
Proposition 3.3.
For any $x\in C_{w}(\bar{K}),h_{\mathcal{L}_{\lambda}}(x)\geq\left\langle\operatorname{deg}F_{Q},w\lambda\right\rangle$.
Proof.
The proof in [3] does not require ${\rm char}(k)=0$, therefore works in our case. ∎
3.2. Height filtration and successive minima
For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\mathcal{L}_{\lambda}}(x)<t\big{\}}$. Let $\zeta_{\operatorname{ess}}(X):=\inf\big{\{}t:Z_{t}=X\big{\}}$ be the essential minimum of $h_{\mathcal{L}_{\lambda}}$ on $X$.
Note that $\mathcal{X}_{w}$ is a closed subscheme of $\mathcal{X}$, so $\mathcal{L}_{\lambda}|_{\mathcal{X}_{w}}$ induces a height function $h_{\mathcal{L}_{\lambda}}:X_{w}(\overline{K})\longrightarrow\mathbb{R}$, which is nothing but the restriction of $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ to $X_{w}(\overline{K})$. Let $\zeta_{\operatorname{ess}}(X_{w})$ be the essential minimum of $h_{\mathcal{L}_{\lambda}}$ on $X_{w}$. Now we compute the essential minimum $\zeta_{\rm ess}(X_{w})$ of $X_{w}$.
Lemma 3.4 ([4], Part I, §5.18).
On $\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P\subseteq\mathcal{X}$, we have
$\pi_{*}\left(\left.\mathcal{L}_{\lambda}\right|_{\mathcal{X}_{w}}\right)=F_{Q}\times_{Q}\mathrm{H}^{0}\left(\overline{QwP}/P,M_{\lambda}\right).$
Let $V$ be any $Q$-representation of highest weight $\lambda\in X^{*}(T)$ and let $V=\bigoplus_{\nu}V[\nu]$ be its weight decomposition. Furthermore, let $F_{Q}$ be the strongly canonical reduction of $F$. We define a filtration $V_{\bullet}$ on the vector space $V$: For any rational number $q\in\mathbb{Q}$, we define the subspace $V_{q}$ as the sum of weight spaces
$V_{q}:=\bigoplus_{\left\langle{\mathrm{deg}}F_{Q},\nu\right\rangle\geq q}V[\nu]$
Clearly, $V_{q^{\prime}}\subseteq V_{q}$ whenever $q^{\prime}\geq q$. We will consider subspaces $V_{q}$ only for the finitely many $q\in\mathbb{Q}$ where a jump occurs, that is, only for those $q$ such that $V_{q^{\prime}}\subsetneq V_{q}$ for all $q^{\prime}>q$. Let $q_{0}$ be the smallest and $q_{1}$ the largest rational number occurring among such $q$. Then $V_{q_{1}}$ is the smallest non-zero filtration step and $V_{q_{0}}$ equals $V$.
Then, by twisting the $Q$-subrepresentations $V_{q}$ above by $F_{P}$, we obtain a filtration $V_{\bullet,F_{Q}}$ of the vector bundle $V_{F_{G}}=F_{G}\times^{G}V$ by subbundles
$0\neq V\left[\lambda+\mathbb{Z}\Delta_{M}\right]_{F_{M}}=(\bigoplus_{\nu\in\lambda+\mathbb{Z}\Delta_{M}}V[\nu])_{F_{M}}=V_{q_{1},F_{Q}}\subsetneq\cdots\subsetneq V_{q,F_{Q}}\subsetneq\cdots\subsetneq V_{q_{0},F_{Q}}=V_{F_{G}}.$
Improving a result of Schieder [^1] The case of characteristic 00 and a specific case of positive characteristic were proven in [8], we proved that:
Proposition 3.5.
Assume the redution $F_{Q}$ is the strongly [^2] Proposition 3.5 fails to hold if we do not require the notion of strongly canonical reduction. canonical reduction, then the filtration $V_{\bullet,F_{Q}}$ of the vector bundle $V_{F_{G}}$ is the Harder-Narasimhan filtration of $V_{F_{G}}$.
We will need the following well-known theorem.
Theorem 3.6 ([6], Theorem 3.23).
If $F$ is a strongly semistable $G$-bundle and $\rho:G\xrightarrow[]{}H$ is a homomorphism that maps the connected component of of the center of $G$ into that of $H$, then $F\times^{G}H$ is a semistable $H$-bundle.
Proof of Proposition 3.5.
It will suffice to check two assertions: all qraded pieces are semistable vector bundles and their slopes are decreasing.
As shown in [8] Proposition 5.1, we have
$\mathrm{gr}_{q}(V_{\bullet,F_{P}})=(\mathrm{gr}_{q}V_{\bullet})_{F_{M}}.$
And its slope is computed as
$\mu((\mathrm{gr}_{q}V_{\bullet})_{F_{M}})=q.$
Therefore, the slopes of the graded pieces are decreasing by definition.
To prove the semistability, consider the semisimplification $W$ of $(\mathrm{gr}_{q}V_{\bullet})$ as a representation of $M$. Now apply [8] Proposition 3.2(a), we see that for each simple direct summand $W_{i}$ of $W$ the associated bundle $W_{i,F_{M}}$ has slope $q$. Thus $(\mathrm{gr}_{q}V_{\bullet})_{F_{M}}$ admits a filtration by subbundles such that each graded piece is a semistable vector bundle of slope $q$, where the semistability is obtained by applying Theorem 3.6 to $M\to{\rm GL}(M_{i})$. Then $(\mathrm{gr}_{q}V_{\bullet})_{F_{M}}$ is also semistable of slope $q$ by the following. ∎
Lemma 3.7.
If $E$ is a vector bundle on $C$ that admits a filtration of subbundles
$0\subsetneq E_{1}\subsetneq\cdots\subsetneq E_{r}=E,$
and all graded piece are semistable of slope equals to $\lambda$, so is $E$.
Corollary 3.8.
Assume $F_{Q}$ is the strongly canonical reduction of $F$. The Harder-Narasimhan filtration of $F_{Q}\times_{Q}H^{0}\left(\overline{QwP}/P,M_{\lambda}\right)$ is $F_{Q}\times_{Q}H^{0}\left(\overline{QwP}/P,M_{\lambda}\right)_{\bullet,\operatorname{deg}\left(F_{Q}\right)}$. Moreover, the maximal slope is $\left\langle\operatorname{deg}\left(F_{Q}\right),w\lambda\right\rangle$.
Proof.
The first assertion is obtained by taking $V$ to be $H^{0}\left(\overline{QwP}/P,M_{\lambda}\right)$ in Proposition 3.5. For the second assertion, it will suffice to show the highest weights in $H^{0}\left(\overline{QwP}/P,M_{\lambda}\right)$ belong to $w\lambda+\mathbb{Z}[\Delta_{Q}]$. The argument in [3] is also valid without any assumption on the characteristic of base field. Note that one can also apply [4, § 14 Proposition 12] in positive characteristic, since Schubert variety in positive characteristic is normal by [7]. ∎
Corollary 3.9.
Under the assumption that $F$ has strongly canonical reduction, the essential minimum $\zeta_{1}(h_{\mathcal{L}_{\lambda}},X_{w})$ of $h_{\mathcal{L}_{\lambda}}$ on $X_{w}$ is $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$.
Proof.
By Ballaÿ’s theorem [1, 2],
$\zeta_{1}(h_{\mathcal{L}_{\lambda}},X_{w})=\lim_{n\to\infty}\frac{\mu_{\rm max}(\pi_{*}\mathcal{L}_{n\lambda}|_{X_{w}})}{n}=\lim_{n\to\infty}\frac{n\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle}{n}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle.$
Proof of Theorem 3.1.
The arguments follows from [3] once the lower bound and essential minimum is obtained. See [3, Proof of Theorem 2.1]
Now we treat the case $F$ may not have a strongly canonical reduction.
Proof of Theorem 3.2.
Theorem 5.1 of [5] shows for a principal $G$-bundle $F$ (might not admits a strongly canonical reduction), the strongly canonical reduction exists for $({\rm Fr}^{n})^{*}F$, when $n$ is sufficiently large. We have the following cartesian diagram
${(\operatorname{Fr_{C}}^{n})^{*}F/P}$${F/P}$${C}$${C}$$\scriptstyle{\phi}$$\scriptstyle{\operatorname{Fr}^{n}}$
with its generic fibre
[…]
Here $\operatorname{Fr_{C}}$ is the absolute Frobenius on $C$. Suppose $n$ is large enough such that $(\operatorname{Fr}^{n})^{*}F$ has strongly canonical reduction. Note that $\phi$ is purely inseparable and we have the equality
$\frac{1}{p^{n}}h_{\phi^{*}\mathcal{L}_{\lambda}}(x)=h_{\mathcal{L}_{\lambda}}(\phi_{K}(x))$
for all $x\in\tilde{X}$. Our theorem follows. ∎